OpenStudy (anonymous):

Jon and Becky measured the circle-shaped part of a sun they drew on the sidewalk. The diameter of the circle is 70 centimeters. What is the approximate area of the circle? Use 3.14 for pi)

OpenStudy (automaticloveletter):

A=π*r^2 sqrt of 70=8.36660027 (to find the radius) A=3.14*8.36^2 A≈219.56

OpenStudy (automaticloveletter):

Medal would be appreciated ^_^